OpenStudy (anonymous):

guys please help me. I give medals A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

To find the time when the particle is first time at the origin, set x(t) = 0. From the nature of cos() function, we know x(t) becomes zero for the first time when t√ equals π2. This gives the time instant. To find the velocity at this time, differentiate x(t) with respect to t to get v(t), and plug in the value of time in the expression for v(t).

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`Find the acceleration` v(t) = s ' (t) a(t) = v ' (t) = s '' (t) so you'll have to find the second derivative of s(t). Equivalently, the first derivative of v(t)

jimthompson5910 (jim_thompson5910):

if s(t) = t*ln(3t), then what is s ' (t) ?

OpenStudy (anonymous):

so the acceleration is the second derivative right?

jimthompson5910 (jim_thompson5910):

yes of the position function

OpenStudy (anonymous):

ok, give me minute

OpenStudy (anonymous):

i think it's wrong, but I got 1+ln(3x) for the first derivative

OpenStudy (freckles):

pretty! Find the acceleration of the particle when the velocity is first zero so you have v=1+ln(3x) you want to find x such that v is 0

OpenStudy (freckles):

then you find v'

jimthompson5910 (jim_thompson5910):

s(t) = t*ln(3t) s ' (t) = d/dt [ t*ln(3t) ] s ' (t) = d/dt [ t ]*ln(3t) + t* d/dt [ ln(3t) ] s ' (t) = 1*ln(3t) + t* (1/3t)*3 s ' (t) = ln(3t) + 1 s ' (t) = 1 + ln(3t) so you have it correct @Joseluess

OpenStudy (anonymous):

ok, so the secod derivative is 1/x

jimthompson5910 (jim_thompson5910):

1/t, yes

OpenStudy (anonymous):

okay so now do I plug zero ?

OpenStudy (freckles):

well it says to evaluate the acceleration when the velocity is 0 that is why I asked you to find x such that v is 0

OpenStudy (freckles):

v=1+ln(3x) 0=1+ln(3x) note: I don't know why it says when the velocity is first 0 because it is only ever 0 once

OpenStudy (anonymous):

but I have to solve for the acceleration, which is the second derivative.

OpenStudy (freckles):

yes I know

OpenStudy (freckles):

but you need to evaluate the acceleration at the value that makes the velocity 0

OpenStudy (anonymous):

so do i solve for x when velocity is zero?

OpenStudy (freckles):

yes solve 0=1+ln(3x)

jimthompson5910 (jim_thompson5910):

`Find the acceleration of the particle when the velocity is first zero` as freckles is saying, you plug v = 0 into v = 1+ln(3t) and solve for t then you take that t value and plug it into a(t) = 1/t

OpenStudy (anonymous):

ok, I got stuck I completely forgot how to solve from this part

OpenStudy (anonymous):



OpenStudy (anonymous):

do I ln both sided to cancel ln ?

OpenStudy (freckles):

\[\log_b(a)=x \text{ has equivalent exponential form } b^x=a \\ \text{ also note that } \ln(x)=\log_e(x)\]

OpenStudy (anonymous):

1/3e

jimthompson5910 (jim_thompson5910):

yes, \[\Large t = \frac{1}{3e}\] now plug that into a(t)

OpenStudy (anonymous):

3e

jimthompson5910 (jim_thompson5910):

yep, a(t) = 3e is your acceleration when t = 1/(3e)

OpenStudy (anonymous):

okay, thank you so much..... I'll give it to you the medal, but @freckles I'll make another question and ill give it to you

OpenStudy (freckles):

no I don't really care I just wanted to help with a calculus problem even though @jim_thompson5910 was already doing an awesome job because lets face it calculus is most fun